Circular No. 7236

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA IN M31 R. Johnson, M. Modjaz and W. D. Li, University of California at Berke-

ley, on behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent nova in M31 (NGC 224). The nova was discovered and confirmed on unfiltered images taken on Aug. 10.4 UT (mag ∼ 15.6) and Aug. 11.4 UT (mag ∼ 16.8). The new object is located at α = 00^{h}42^{m}22^{s}.1, δ = +41^{o}13′23′′ (equinox 2000.0). A KAIT image of the same field on Aug. 9.4 UT (limiting mag ∼ 19.0) showed nothing at the position of the nova. A finding chart can be found at

http://astron.berkeley.edu/~bait/1999/n31 2.html.

V382 VELORUM Visual magnitude estimates: 1999 July 21.972 UT, 7.6 (A. Amorim,

Florianopolis, Brazil); 24.956, 7.8 (J. Hodar, Campinas, Brazil); 25.054, 7.7 (J. Garcia, Rama Caida, Brazil); 26.910, 7.7 (T. A. Napoleao, São Paulo, Brazil); 26.987, 7.7 (R. Salvo, Montevideo, Uruguay); 29.899, 7.8 (J. G. de S. Aguiar, Campinas, Brazil); 29.937, 7.7 (W. C. de Souza, Pico dos Marins, Brazil, 7 × 50 binoculars); 31.837, 7.8 (A. Alves, Florianopolis, Brazil); 31.934, 7.7 (M. Mendes, Dois Corregos, Brazil); Aug. 1.890, 7.8 (Aguiar); 3.885, 7.9 (Aguiar); 6.888, 7.9 (Aguiar); 10.889, 8.0 (Aguiar).

R CORONAE BOREALIS Visual magnitude estimates: 1999 July 5.96 UT, 6.2 (J. D. Shanklin,

Cambridge, England); 17.96, 6.4 (Shanklin); 23.8556, 6.1 (S. Foglia, Milan, Italy, 10 × 50 binoculars); 24.8535, 6.4 (Foglia); 31.97, 6.7 (A. Pereira, Cabo da Roca, Portugal); Aug. 1.908, 6.8 (K. Hornoch, Lelekovice, Czech Republic, 10 × 80 binoculars); 6.838, 7.1 (Hornoch); 10.8743, 7.5 (Foglia); 10.98, 7.7 (Shanklin); 11.8431, 7.6 (Foglia).

CATALOGUE OF COMETARY ORBITS 1999 Further to the item on IAUC 7235, orders for the 1999 edition of the

Catalogue of Cometary Orbits may be sent to iausubs@cfa.harvard.edu or to the FAX number or postal address at the top of this Circular. Checks should be made out to ‘Central Bureau for Astronomical Telegrams’. Payment can be made by credit card. Subscribers to these Circulars can have their accounts debited by the appropriate amount.

1999 August 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Gareth V. Williams